OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? x2 - 7x - 8 A.x - 4 B.x + 2 C.x + 4 D.x - 8

OpenStudy (tkhunny):

Have you considered factoring the Trinomial? x^2 - 7x - 8 = (x + ___)(x - ___) Fill in the blanks with factors of 8. 8 = 1 * 8 8 = 2 * 4

OpenStudy (anonymous):

so would the answer be x+4? @tkhunny

OpenStudy (tkhunny):

You tell me. Did you factor it?

OpenStudy (tkhunny):

Let's see. 4 * 8 = 32. Nope! That is not a proper factorization. x^2 - 7x - 8 = (x + ___)(x - ___) Fill in the blanks with factors of 8. 8 = 1 * 8 8 = 2 * 4 It's either +1 and -8 or -1 and +8 or -2 and +4 or +2 and -4. No other possibilities.

OpenStudy (anonymous):

wait would it be (x+1)(x-8)? so it would be D? @tkhunny

OpenStudy (tkhunny):

Perfect. Now the question is simple.